OpenStudy (anonymous):

using the multiplication principle solve -8c=-24

OpenStudy (anonymous):

c=3

OpenStudy (anonymous):

how do we get answer

OpenStudy (anonymous):

c = -24/-8 c = 3

OpenStudy (anonymous):

for example, this: 2*3=6 to find 3, 3 = 6/2 3 is here as variable "c" in your question

OpenStudy (anonymous):

i recommend you to use this method, modelling method

OpenStudy (anonymous):

would that be a -3

OpenStudy (anonymous):

no, because if you divide negative to negative number you get positive number memorize it!

OpenStudy (anonymous):

-24 and -8 are both negative numbers for example \[(-4) \times (-5) = 20\]

OpenStudy (anonymous):

oh ok got it thanks

OpenStudy (anonymous):

glad to help!